OpenStudy (anonymous):

If cosine of x equals 1 over 2, what is sin(x) and tan(x)? Explain your steps in complete sentences.

OpenStudy (freckles):

make a right triangle then use cos(x)=1/2 then use Pythagorean theorem to figure out opposite side of x

OpenStudy (freckles):

then you have everything you need to find sin(x) and tan(x)

OpenStudy (anonymous):

can u guide me @freckles

OpenStudy (freckles):

which part do you need help on?

OpenStudy (freckles):

did you draw the right triangle yet?

OpenStudy (anonymous):

i guess all.

OpenStudy (anonymous):

yes

OpenStudy (freckles):

ok then use cos(x)=1/2 remember cosine, think adjacent/hyp

OpenStudy (anonymous):

ok

OpenStudy (freckles):

then use the Pythagorean theorem to find the opp side can you do that?

OpenStudy (anonymous):

no

OpenStudy (freckles):

well we know the Pythagorean theorem says: \[opp^2+adj^2=hyp^2 \\ \text{ and we know the } adj \text{ and the hyp } \\ \text{ just solve for opp}\]

OpenStudy (anonymous):

adj is 1 and hyp is 2?

OpenStudy (freckles):

yes this was given by cos(x)=1/2 you could also use any thing that reduces down to 1/2 for example cos(x)=10/20 and call the adj=10 and hyp=20 but it doesn't matter in the end we will get the same reduced fractions so yes you can use adj=1 and hyp=2

OpenStudy (anonymous):

ok

OpenStudy (freckles):

have you solved for opp yet?

OpenStudy (freckles):

if you have you can find sin(x) and tan(x)

OpenStudy (anonymous):

still working on it

OpenStudy (anonymous):

opp is root 5?

OpenStudy (freckles):

\[oop^2+1^2=2^2 \\ oop^2+1=4\] 4-1 isn't 5 should be 3

OpenStudy (freckles):

\[opp^2=4-1 \\ opp^2=3 \\ opp= \pm \sqrt{3}\]

OpenStudy (anonymous):

shoot my bad

OpenStudy (freckles):

so you have two answers for both sin(x) and tan(x) because of the plus or minus thing

OpenStudy (freckles):

anyways recall sin(x)=opp/hyp and tan(x)=opp/adj

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sin(x) = 3/2 tan(x) = 3/1 = 3

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

what happen to the sqrt( ) part ?

OpenStudy (freckles):

also you should have two answers for each

OpenStudy (freckles):

\[\sin(x)=\frac{ \pm \sqrt{3}}{2} \text{ this is the two answers for } \sin(x) \] now you try tan(x)=?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

±root 3?

OpenStudy (freckles):

yep sounds great

OpenStudy (anonymous):

how did you get the root sighn tho

OpenStudy (anonymous):

sign

OpenStudy (anonymous):

jk never mind

OpenStudy (freckles):

\[opp^2+adj^2=hyp^2 \\ opp^2+1^2=2^2 \\ opp^2=4-1 \\ opp^2=3 \\ \text{ \to solve an equation like this you take square root of both sides }\]

OpenStudy (freckles):

but it will also mean you wind up with two possibilities since sqrt(opp^2)=|opp|=opp if opp>0 but =-opp if opp<0

OpenStudy (anonymous):

ok. Thank You so so so so much!!!!!!!!!!!!!!!!!!!